# Hardwired-Neurons Language Processing Units as General-Purpose Cognitive Substrates

### v1

@@ Abstract.
The rapid advancement of Large Language Models (LLMs) has established language as a core general-purpose cognitive substrate, driving the demand for specialized Language Processing Units (LPUs) tailored for LLM inference. To overcome the growing energy consumption of LLM inference systems, this paper proposes a Hardwired-Neurons Language Processing Unit (HNLPU), which physically hardwires LLM weight parameters into the computational fabric, achieving several orders of magnitude computational efficiency improvement by extreme specialization. However, a significant challenge still lies in the scale of modern LLMs. An ideal estimation on hardwiring gpt-oss 120 B requires fabricating at least 6 billion dollars of photomask sets, rendering the straightforward solution economically impractical.
Addressing this challenge, we propose the novel Metal-Embedding methodology. Instead of embedding weights in a 2D grid of silicon device cells, Metal-Embedding embeds weight parameters into the 3D topology of metal wires. This brings two benefits: (1) a 15× increase in density, and (2) 60 out of 70 layers of photomasks are made homogeneous across chips, including all EUV photomasks. In total, Metal-Embedding reduced the photomask cost by 112×, bringing the Non-Recurring Engineering (NRE) cost of HNLPU into an economically viable range. Experimental results show that HNLPU achieved 249,960 tokens/s (5,555×/85× of GPU/WSE), 36 tokens/J (1,047×/283× of GPU/WSE), 13,232 mm2 total die area (29% inscribed rectangular area in a 300 mm wafer), \$ 184 M estimated NRE at 5 nm technology. Analysis shows that HNLPU achieved 8.57× cost-effectiveness and 230× carbon footprint reduction compared to H100 clusters, under an annual weight updating assumption.
@@ 1. Introduction
The current trends in artificial intelligence development indicate that language capabilities serve as the core foundation for constructing advanced cognitive and learning systems. This is exemplified by the rapid advancement of Large Language Models (LLMs), which have achieved technological unification in Natural Language Processing (NLP) (Vaswani et al., 2017; Brown et al., 2020; Wei et al., 2022). By demonstrating human-level performance across diverse downstream tasks, LLMs are expanding their application scope while exhibiting a growing trend toward enhanced generality—enabling a single model to handle multiple tasks that previously required distinct systems.
@@
Driven by the widespread adoption while heavy hardware overhead of LLMs, there has been a surge in demand for dedicated processors specifically designed for LLM inference, collectively known as Language Processing Units (LPUs). For instance, Groq LPU(Abts et al., 2020, 2022) and Cerebras WSE(Lauterbach, 2021; Lie, 2023) pre-load model weights into on-chip SRAM, while Etched Sohu(etc, 2025) hardens the transformer dataflow into its compute fabric. These LPUs leverage the advanced specialization targeting LLMs, achieving $4\sim 20\times$ energy efficiency compared to traditional CPUs, GPUs, and NPUs.
@@
Figure 1. Hardwired-Neurons LPU as a general-purpose processor. To the left: AI Infrastructures are originated in the rapidly evolving deep learning which appreciates universality over extreme efficiency. To the right: As LLM develops, the responsibilities of universality are shifting from HW/SW to LLMs. An extremely specialized Hardwired-Neurons LPU can also be helpful in general tasks.
@@
Although LPUs have brought a certain improvement in hardware efficiency, AI infra still suffers from tremendous energy cost. For example, AI datacenters are estimated to occupy 12% of the total electricity capacity in US before 2028, which is clearly unsustainable (of Energy, 2024). This is due to the fact that current LPUs and GPUs still see models as dynamically changing data, despite the fact that leading enterprises typically deploy only a single or a few proprietary LLM. The hundreds of billions of weight parameters are repeatedly fetched during each autoregressive decoding step, consuming the majority of system power.
@@
To fundamentally solve the limitation, we argue to push the specialization of LPU to the extreme. By physically hardwiring weight parameters into the LPU fabric, Hardwired-Neurons LPU (HNLPU) can achieve perfect architecture-model matching, zero parameter fetching overhead, and extreme computational efficiency from constants-arithmetic circuits. As one dominant pre-trained LLM can serve as a general-purpose cognitive substrate for a wide variety of tasks, the extreme specialization once considered too inflexible is transforming into a reasonable choice with huge potential benefits (Figure 1).
@@
However, previous attempts on HNLPU would be stopped by the extraordinarily large scale of modern LLM. The most optimistic estimation on hardwiring gpt-oss 120 B (OpenAI, 2025) would require 176,000 mm2 multiply-by-constant units (CMAC) array at 5 nm technology. Unlike previous wafer-scale practices stepping a repeated lithographic pattern on the wafer, here the patterns are heterogeneous everywhere, because embedded weight parameters are different everywhere. The photomask sets valued over 6 billion dollars, rendering the straightforward implementation of HNLPU economically prohibitive.
Figure 2. Conceptual schematic of Metal-embedding. ❶ Conventional design embeds weights with 2D grid of cells. Metal layers are only for physical design (P&R). ❷ Metal-embedding dramatically increases the density by leveraging the 3D topology structure of metal wires, emulating the axon-dendrite in brains.
@@
Economically viable HNLPU must come after a weight-embedding methodology breakthrough. In this paper, we propose the Metal-Embedding (ME) methodology, achieving multiple order-of-magnitudes reduction on photomask counts required to implement HNLPU. Roughly speaking, instead of embedding weights in the 2D grid of cells (either MAC/CMAC/SRAM/ReRAM cells), ME embeds weights in the 3D topology of metal wires (Figure 2). The benefits of ME is two-fold. First, due to the inherently richer expression of 3D structure, ME reduced the area by an order-of-magnitude (-93.4%) compared with CMAC grid. Second, these metal wires can be placed at higher metal layers (M8+), to be cost-efficiently fabricated with trailing-edge optical lithography (193i). As a result, ME keeps photomask homogeneous (shared) across all chips for all layers in Front-End-of-Line (FEOL) and critical layers in Back-End-of-Line (BEOL), including all Extreme Ultraviolet (EUV) photomask in the process. This reduced the total cost on photomask by another order of magnitude: -86.5% for initial tapeout, -92.3% for parameter-only update re-spin.
@@
We evaluate HNLPU at 5 nm technology based on post-layout characteristics, and compare with baselines (NVIDIA H100 GPU, Cerebras WSE-3) at the same technology node. Experimental results demonstrate that HNLPU achieved 249,960 tokens/s throughput (5,555× of H100, 85× of WSE-3), 36 tokens/J energy efficiency (1,047× of H100, 283× of WSE-3), at 13,232 mm2 total die area divided into 16 chips. Estimated Non-Recurring Engineering cost (NRE) of HNLPU is \$ 184 M for the initial tapeout, \$ 44.3 M for a parameter-only update re-spin. Throughput per 3-Year Total Cost of Ownership (TCO) is 8.57×/12.65× of H100 clusters with/without annual parameter updates, and the total carbon footprint (tCO2e) is reduced by 230×/234×.
@@
  • We explored the concept and feasibility of HNLPU for the first time.
  • We propose the Metal-Embedding methodology, reduced the photomask cost of HNLPU by two orders of magnitude.
  • We detailed the architecture and dataflow of the first HNLPU design. It implemented a gpt-oss 120 B (FP4) in 13,232 mm2 total die area, achieving unprecedented throughput and energy efficiency.
@@ 2. Background and Motivation
2.1. Language as a cognitive substrate
Current AI development trends underscore language as the foundational substrate for constructing advanced cognitive capabilities. This centrality is supported by three key arguments. First, language is demonstrably critical to human intelligence. Neurocognitive research reveals strong correlations between linguistic competence and higher-order cognitive abilities, such as abstract reasoning (Lupyan and Clark, 2015). These findings strongly suggest that language should similarly form the core of advanced artificial intelligence. Second, the sheer abundance of textual data vastly exceeds that available in other modalities. For instance, Common Crawl (Common Crawl, [n. d.]) has amassed an 8 PB text corpus from web pages, growing consistently by approximately 250 TB per month. This unparalleled scale of pre-training data is instrumental in enabling the zero-shot generalization abilities observed in Large Language Models (LLMs). Third, language constitutes the most natural and universal interface for human-AI collaboration. Robust language understanding and generation capabilities are indispensable for AI to integrate into human workflows and societal structures.
The rise of LLMs has further cemented language’s foundational status through technical unification. Historically, NLP was fragmented into specialized subfields (e.g., sentiment analysis, machine translation, named entity recognition), each requiring task-specific systems. Today, a single pre-trained LLM achieves state-of-the-art performance across common NLP tasks. This convergence stems from scaling model parameters and training data, enabling one general-purpose model to outperform prior specialized approaches. Consequently, LLMs now serve as universal engines for language-powered cognition, spanning reasoning, code generation, and multimodal interaction through linguistic interfaces.
This unification demonstrates that a single, highly capable LLM can function as a general-purpose cognitive substrate. While multiple implementations exist (e.g., GPT, Claude, Gemini, Qwen and DeepSeek), the focus of competition is shifting from capabilities to computational efficiency, a domain where model-architecture co-design becomes decisive. The demonstrated generality of modern LLMs makes architectural specialization strategically viable. For instance, models like gpt-oss 120 B which achieves state-of-the-art performance, exemplify how a singular, optimally scaled Transformer architecture can serve as a universal language engine. By fixing such a model as a standardized substrate, unprecedented hardware optimizations become feasible, while maintaining full generality through the model’s inherent language-based cognitive flexibility.
2.2. Specialized Language Processing Units
The emergence of LLM as universal cognitive engines has intensified demand for dedicated inference hardware. We define processors explicitly optimized for language models as Language Processing Units (LPUs), a complementary paradigm following CPUs/GPUs/NPUs. However, currently LPUs are predominantly repurposed from general-purpose processors. For example, the most widely-deployed NVIDIA Hopper architecture incorporates dedicated Transformer Engines, but still retains the architectural lineage of its Tensor Core GPUs originated before LLMs. These repurposed processors suffer from efficiency limitations intrinsic to their origin in more general applications, known as ”Turing Tariffs” (Kelly, 2020).
The major efficiency limitation in current architectures is two-fold. First, the architecture-model mismatch fails to accommodate divergent computational demands across inference phases (e.g., compute-bound prefill vs. memory-bound decoding), forcing industrial compromises like prefill/decode and attention/feed-forward decoupling that deploy heterogeneous hardware subsystems specifically (Aminabadi et al., 2022; Wang et al., 2025). Second, repurposed processors dynamically load models as ever-changing data, despite the fact that enterprises typically deploy only a single or a few proprietary LLM. The >>100 B model parameters are repeatedly fetched during each autoregressive decoding step, dominating the whole system power.
Emerging architectures like Cerebras Wafer-Scale Engine (WSE) and Groq LPU attempt to circumvent these bottlenecks through deterministic scheduling and massive on-chip SRAM integration, reportedly achieved up to 20× energy efficiency gain over GPUs (Abts et al., 2020; Lie, 2023). These practices demonstrated the great benefits and viability of pushing the specialization of LPU hardware. However, Turing Tariffs still exist in these architectures. Only on one hand, parameter fetching still dominates the power consumption as the SRAM access cost is typically several times higher than computation (Sze et al., 2017).
To fundamentally address these limitations, we propose to push the specialization of LPU to an architectural extreme: permanent weight hardwiring within LPU chips. By physically embedding model parameters into the fabric of computing units, $\sim$1000× efficiency gain could be obtained from perfect architectural matching, zero parameter fetching overhead, and constant arithmetic. Such model-specialized architecture aligns with the industrial consolidation trend, where enterprises increasingly standardize on one dominant LLM architecture, transforming what was once considered an inflexibility into a fairly reasonable choice with huge potential benefits.
2.3. Hardwiring a Language Processing Unit
@@
A glaring challenge has thwarted all ambitious attempts to hardwire the LPUs: LLMs are too large for the current lithographic technology. For example, the typical transistor density of high-density 5 nm technology is around 138 MTr/mm2. FP4 Constant MAC (CMAC) requires 200+ transistors. This translates gpt-oss 120 B into an ideal area estimation of 176,000 mm2, four maximal inscribed rectangles in 300 mm wafer. Due to reticle size limits, this hardwired LPU has to split into 200+ sets of photomask. Unlike prior wafer-scale practices, these masks are heterogeneous. This incurs 6 billion dollars non-recurring engineering (NRE) costs on photomask making, rendering a straightforward hardwiring LPU economically prohibitive.
@@
Compounding this economic infeasibility is the fast LLM development cycles. The absurdly high NRE is even worse when considering the annual parameter updates and 3-year typical lifespan of production LLMs (e.g., GPT-4). This multiple-orders-of-magnitude cost gap is unlikely to be bridged solely through model compression. Please note that the gpt-oss is already FP4 by default. The model size has a concrete lower bound implied by the Kolmogorov complexity of human knowledge representation.
@@
To realize the concept of hardwired LPUs, there must be fundamental breakthroughs in the on-die weight embedding methodologies. By co-optimizing with lithography technology factors, we can achieve multiple orders-of-magnitudes reduction in photomask count, eventually clear the path to the economically viable hardwired LPUs.
3. Design Principles

### v2

@@ Abstract.
The rapid advancement of Large Language Models (LLMs) has established language as a core general-purpose cognitive substrate, driving the demand for specialized Language Processing Units (LPUs) tailored for LLM inference. To overcome the growing energy consumption of LLM inference systems, this paper proposes a Hardwired-Neurons Language Processing Unit (HNLPU), which physically hardwires LLM weight parameters into the computational fabric, achieving several orders of magnitude computational efficiency improvement by extreme specialization. However, a significant challenge still lies in the scale of modern LLMs. A straightforward hardwiring of gpt-oss 120 B would require fabricating photomask sets valued at over 6 billion dollars, rendering this straightforward solution economically impractical.
Addressing this challenge, we propose the novel Metal-Embedding methodology. Instead of embedding weights in a 2D grid of silicon device cells, Metal-Embedding embeds weight parameters into the 3D topology of metal wires. This brings two benefits: (1) a 15× increase in density, and (2) 60 out of 70 photomask layers are homogeneous across chips, including all EUV photomasks. In total, Metal-Embedding reduced the photomask cost by 112×, bringing the Non-Recurring Engineering (NRE) cost of HNLPU into an economically viable range. Experimental results show that HNLPU achieved 249,960 tokens/s (5,555×/85× that of GPU/WSE), 36 tokens/J (1,047×/283× that of GPU/WSE), 13,232 mm2 total die area, \$ 59.46 M–123.5 M estimated NRE at 5 nm technology. Analysis shows that HNLPU achieved 41.7–80.4× improvement in cost-effectiveness and 357× reduction in carbon footprint compared to OpenAI-scale H100 clusters, under an annual weight updating assumption.
†Corresponding Author: Qi Guo
*Contribute Equally
@@ 1. Introduction
Current trends in artificial intelligence development indicate that language capabilities serve as the core foundation for constructing advanced cognitive and learning systems. This is exemplified by the rapid advancement of Large Language Models (LLMs), which have achieved technological unification in Natural Language Processing (NLP) (Vaswani et al., 2017; Brown et al., 2020; Wei et al., 2022). By demonstrating human-level performance across diverse downstream tasks, LLMs are expanding their application scope while exhibiting a growing trend toward enhanced generality—enabling a single model to handle multiple tasks that previously required distinct systems.
@@
Driven by widespread adoption despite the heavy hardware overhead of LLMs, there has been a surge in demand for dedicated processors specifically designed for LLM inference, collectively known as Language Processing Units (LPUs). For instance, Groq LPU (Abts et al., 2020, 2022) and Cerebras WSE (Lauterbach, 2021; Lie, 2023) pre-load model weights into on-chip SRAM, while Etched Sohu (Wikipedia contributors, 2025) hardens the transformer dataflow into its compute fabric. These LPUs leverage the advanced specialization targeting LLMs, achieving $4\sim 20\times$ energy efficiency compared to traditional CPUs, GPUs, and NPUs.
@@
Figure 1. Hardwired LPU as a general-purpose processor. To the left: AI Infrastructures are originated in the rapidly evolving deep learning which appreciates universality over extreme efficiency. To the right: As LLM develops, the responsibilities of universality are shifting from HW/SW to LLMs. An extremely specialized Hardwired LPU can also be helpful in general tasks.
@@
Although LPUs have brought a certain improvement in hardware efficiency, AI infrastructure still suffers from tremendous energy costs. For example, AI datacenters are estimated to occupy 12% of the total electricity capacity in the US by 2028, which is clearly unsustainable (of Energy, 2024). This is due to the fact that current LPUs and GPUs still see models as dynamically changing data, despite the fact that leading enterprises typically deploy only a single or a few proprietary LLMs. The hundreds of billions of weight parameters are repeatedly fetched during each autoregressive decoding step, consuming the majority of system power.
@@
To fundamentally solve the limitation, we argue for pushing the specialization of LPU to the extreme. By physically hardwiring weight parameters into the LPU fabric, Hardwired LPU can achieve perfect architecture-model matching, zero parameter fetching overhead, and extreme computational efficiency from constants-arithmetic circuits. As one dominant pre-trained LLM can serve as a general-purpose cognitive substrate for a wide variety of tasks, the extreme specialization once considered too inflexible is transforming into a reasonable choice with huge potential benefits (Figure 1).
@@
However, previous attempts on hardwiring would be stopped by the extraordinarily large scale of modern LLM. The most optimistic estimation on hardwiring gpt-oss 120 B (OpenAI, 2025) would require a 176,000 mm2 constant-multiply-and-accumulate units (CMAC) array at 5 nm technology. Unlike previous wafer-scale practices that step a repeated lithographic pattern on the wafer, here the patterns are heterogeneous everywhere, because embedded weight parameters are different everywhere. The photomask sets are valued at over 6 billion dollars, rendering the straightforward implementation of Hardwired LPU economically prohibitive.
@@
An economically viable Hardwired LPU requires a breakthrough in weight-embedding methodology. In this paper, we propose the Metal-Embedding (ME) methodology, achieving multiple orders-of-magnitude reduction on photomask counts required to implement Hardwired LPU. Roughly speaking, instead of embedding weights in the 2D grid of cells (either MAC/CMAC/SRAM/ReRAM cells), ME embeds weights in the 3D topology of metal wires (Figure 6). The benefits of ME are two-fold. First, due to the inherently richer expression of 3D structure, ME reduced the area by an order-of-magnitude (-93.4%) compared with the CMAC grid. Second, these metal wires can be placed at higher metal layers (M8+), to be cost-efficiently fabricated with trailing-edge optical lithography (193i). As a result, ME keeps photomask homogeneous (shared) across all chips for all layers in Front-End-of-Line (FEOL) and critical layers in Back-End-of-Line (BEOL), including all Extreme Ultraviolet (EUV) photomask in the process. This reduced the total cost on photomask by another order of magnitude: -86.5% for initial tapeout, -92.3% for parameter-only update re-spin.
@@
We evaluate the first Hardwired LPU design HNLPU at 5 nm technology based on post-layout characteristics, and compare with baselines (NVIDIA H100 GPU, Cerebras WSE-3) at the same technology node. Experimental results demonstrate that HNLPU achieved 249,960 tokens/s throughput (5,555× of H100, 85× of WSE-3), 36 tokens/J energy efficiency (1,047× of H100, 283× of WSE-3), at 13,232 mm2 total die area divided into 16 chips. Estimated Non-Recurring Engineering cost (NRE) of HNLPU is \$ 59.46 M–123.5 M for the initial tapeout, \$ 18.53 M–37.06 M for a parameter-only update re-spin. Under the OpenAI-scale deployment, the 3-Year Total Cost of Ownership (TCO) is lowered by 41.7–80.4× compared to H100 clusters with annual parameter updates under the OpenAI-scale volume, and the total carbon footprint (tCO2e) is lowered by 357×.
@@
  • We explored the concept and feasibility of Hardwired LPU for the first time.
  • We propose the Metal-Embedding methodology, reduced the photomask cost of Hardwired LPU by two orders of magnitude.
  • We detailed the architecture and dataflow of the first Hardwired LPU design, HNLPU. It implemented a gpt-oss 120 B (FP4) in 13,232 mm2 total die area, achieving unprecedented throughput and energy efficiency.
@@ 2. Background and Motivation
2.1. Hardwired Language Processing Units
In principle, orders-of-magnitude gains in computational efficiency can be achieved by directly fabricating neural network models into hardware. Since the 1980s, hardwired neural networks have been implemented in VLSI (Graf et al., 1988), optical (Abu-Mostafa and Psaltis, 1987; Xu et al., 2024), and printed flexible (Mubarik et al., 2020; Weller et al., 2021; Ozer et al., 2023, 2020; Chakraborty, 2024) circuits. However, the diversity and rapid evolution of neural networks have rendered such hardwired implementations of limited practical value and prone to rapid obsolescence. As a result, instead of hardwired neural networks, general-purpose processors such as GPUs and NPUs have prevailed due to their programmability. Contemporary AI infrastructure and software stacks including programming languages (e.g., CUDA, Triton), libraries (e.g., cuDNN, flash-attn), frameworks (e.g., PyTorch, vLLM) are built on a sweet spot between computational efficiency and architectural generality.
The rise of large language models (LLMs) has fundamentally altered this landscape. Unlike earlier task-specific neural networks designed for isolated tasks, LLMs function as general-purpose cognitive substrates capable of reasoning, planning, and interacting across a wide range of domains. For the first time in history, we possess a single neural network model that is valuable enough to justify hardwired implementation.
Although technically hardwired (in the sense that its parameter weights are physically immutable), the circuit embodies a new paradigm of the general-purpose processor. We refer to this novel processor paradigm as the Hardwired Language Processing Unit (Hardwired LPU). By leveraging the emergent capabilities of in-context learning and zero-shot reasoning, a hardwired LPU treats natural language prompts as a high-level instruction stream, effectively replacing the traditional binary Instruction Set Architecture (ISA) with a semantic interface. Consequently, the user programs the Hardwired LPU not by altering its models and weights, but by prompting with tokens to perform arbitrary tasks.
In this paper, we demonstrate that a single-node Hardwired LPU can outperform a middle-sized GPU cluster without requiring expensive software support, thus establishing a decisive operational-expenditure (OpEx) advantage over challengers. Therefore, we envision Hardwired LPUs as game-changing platforms for long-term, high-volume deployment, while GPU clusters assume the role of short-cycle model development and evaluation testbeds.
2.2. Economic Challenges
Figure 2. Economic challenges of hardwiring. Considering the cost on photomasks and wafers, the cost on photomasks was amortized by the mass production of GPUs. Hardwiring an LLM incurs too many photomasks and too low volume to amortize the NRE costs.
A glaring challenge has thwarted all ambitious attempts to hardwire LLMs: LLMs are too large for the current lithographic technology. The practical obstacle lies in the non-recurring engineering (NRE) cost, primarily the cost from photolithographic masks.
In semiconductor technology, chips are fabricated as layered structures defined by photomasks (analogous to stencils). Front-end-of-line (FEOL) processes, which form the device cells (transistors) on the silicon wafer, typically require on the order of 30 mask layers, while back-end-of-line (BEOL) processes, which form the metal interconnect stack, require approximately 40 additional mask layers. At 5 nm technology nodes, a full photomask set is valued at \$ 30 M, while each processed wafer is valued at approximately \$ 18,000. As shown in Figure 2, the high NRE cost of photomasks is meant to be amortized with massive production. For example, NVIDIA is estimated manufacturing 20,000 wafers of H100 GPU, then the amortized photomask cost would be \$ 1,500 per wafer. However, the expected production volume of hardwired LPUs is very low. Due to the extreme computational efficiency, tens of wafers would oversaturate current LLM service demand on the planet, which renders the expensive photomask set for few uses only.
@@ 2.2. Economic Challenges
A straightforward implementation would require an extraordinarily large silicon area to embed the weights, and a correspondingly large number of photomask sets. For example, the typical transistor density of high-density 5 nm technology is around 138 MTr/mm2. FP4 Constant MAC (CMAC) requires 200+ transistors. This translates gpt-oss 120 B into an ideal area estimation of 176,000 mm2 divided into 200+ chips. Due to reticle size limits, this hardwired LPU must be split across 200+ photomask sets. Unlike prior wafer-scale practices, these masks are heterogeneous because each chip is carrying different parts of the model weights. This incurs \$ 30 M $\times$ 200 = \$ 6 B NRE costs on photomask making, rendering a straightforward hardwired LPU economically prohibitive.
@@
Compounding this economic infeasibility is the fast LLM development cycles. The prohibitively high NRE is even worse when considering the annual parameter updates and 3-year typical lifespan of production LLMs (e.g., GPT-4). This multiple-orders-of-magnitude cost gap is unlikely to be bridged solely through model compression. Note that the gpt-oss is already FP4 by default. The model size has a concrete lower bound implied by the Kolmogorov complexity of human knowledge representation.
@@
To realize the concept of hardwired LPUs, there must be fundamental breakthroughs in on-die weight embedding methodologies. By co-optimizing with lithographic technology factors, we can achieve multiple orders-of-magnitude reduction in photomask count, eventually clear the path to the economically viable hardwired LPUs.
3. Metal-Embedding
We address this economical challenge with the novel Metal-Embedding (ME) methodology. There are two key innovations in ME. 1) The Hardwired-Neuron (HN) architecture rearranges conventional multiply-accumulate arithmetic units into accumulate-multiply-accumulate, and lifted the embedding of weight parameters from silicon devices into metal interconnections. This enables 2) the Sea-of-Neurons architecture – a metal-programmable structured ASIC saving photomasks through a prefabricated array of HNs.
3.1. Hardwired-Neuron Architecture
We demonstrate the step-by-step evolution from the straightforward FP4 multiply-and-accumulate units to Hardwired-Neurons (HN). Several key arithmetic techniques are applied to minimize the required silicon area.
Basic: Weight Constancy
Conventional hardwiring (the \$ 6 B scenario) utilizes weight constancy. By fixing the weights, multipliers could be optimized as multiply-by-constant which is several times lower in Boolean complexity. An FP4 multiply-by-constant unit is $\sim$6× smaller than an FP4 multiplier as seen in GPU. Accumulation could also benefit from the weight constancy under the help of optimizing EDA tools.
Figure 3. Key arithmetic techniques. To the middle: Combining repeated multipliers via the distributive law. To the right: Using Carry Save Adders (CSA) on bit-serialized inputs to trade time for area.
Step 1: Distributive Law
In conventional hardwiring, FP4 weights have 16 unique values, but there are 2,880111The hidden size in gpt-oss 120 B. constant multipliers in each neuron. Most of them are repeated. By the distributive law, common multipliers could be extracted and combined. As shown in Figure 3, instead of performing $ax_{1}+ax_{2}+\dots+ax_{n}$ (to the left), HN performs $a\left(x_{1}+x_{2}+\dots+x_{n}\right)$ (to the middle) which saves multipliers and reduces the width of accumulation.